<$BlogRSDURL$>

4.05.2004


3.31.2004

2004-0-31 12:12pm RCB

Went straight here to IDC had rough night last night, game was still erratic but needed to catch some Zs. Tito and Marc already here, before coming over talked to Tito over the phone and agreed we should try reverting back to de-aggregate the router/switch link. True enough guys noticed that the lights were turning on and off, the connections from the servers to the aggregate swicth. Have re-configured it and are now monitoring. Keeping fingers cross.

3.30.2004

2004-03-30 6:26pm

things were still erratic after the line went up. items were missing, denied log-ins, sexes of avaters are being changed and our E1 link connection to the servers and even email are on the fritz. we did test to ping the servers from the outside and we cannot reach them. internally however, we are able to access them i.e. the web server. tito reset the firewall for the common servers and the web became available.

that wasnt the end of that. link is still erratic and gravity noticed that the clocks of the servers are not synchronized. our presumption is that it all got screwed up when pldt did the reset of the router and that trickled down the servers.

clocks were re-synched, servers / databases were reset as well as the network devices. things look a little better, there are still some items missing and the characters in Iris are not showing up. The characters however are still there according to the gm tool. waiting for other problems to materialize (or hopefully dissappear)
2004-03-30 3:14pm

Link established around 1:30pm.
Problem seemed to be from GBIC Card of the Cisco router. However, people are unable to connect to the servers from the outside (external). Spare GBIC Card will be delivered by Trends this 4:30pm.
Tito went to RC and reset the firewall and it seemed to be fixed.



2004-03-30 (Tuesday)

Around 7:00am Jaq sent me an SMS message that the Fiber link from RCB to Vitro bogged down. Consequently sent SMS of same to Ben and Gil.

Left for office and SMS again Jaq. Was told the Igate people had gone there to investigate and waited for further inputs.

Traffic bad, reached office around 9:30am Mark Chan informed me that people are still investigating. SMS Joseph Versoza and was told that restoration is still ongoing. Told him to keep me posted.SMS the same to Ben and Gil. Its now 9:50am still awaiting update.

This page is powered by Blogger. Isn't yours?